The measures of the three angles of a triangle has the ratio of 3:4:5. What are the measurement of the angles?
Note: Please include representations of any used variable. Say, Let x be.....
Thank you very much..

Answer by checkley77(12844)

3+4+5=12
90/12=7.5
3*7.5=22.5 deg. is the smaller angle.
4*7.5=30 deg. is the middle angle.
5*7.5=37.5 deg. is the larger angle.
Proof:
22.5+30+37.5=90
90=90